OpenStudy (cedric):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (cedric):

I believe its C due to the fact the first to dont work and D is implausiblle

OpenStudy (anonymous):

It D

OpenStudy (cedric):

How?

OpenStudy (anonymous):

They have the same slope

OpenStudy (anonymous):

First, determine if the slopes of the equations are the same. If yes, then it would be no solution or infinitely many solution. For no solution, they are parallel line. They have the same slope but different y-int (or x-int) For infinitely many solutions, they have the same slope and same y-int/x-int. In this case, I think @cedric you're right!

OpenStudy (anonymous):

sorry i didn't see the negative sign my bad

OpenStudy (cedric):

one more?

OpenStudy (anonymous):

Even if there is no negative sign, it's not D, as the intercepts are different!

OpenStudy (anonymous):

2x + y = 1 4x + 2y = −1 Get y by itself on 2x+y=1 -2x -2x y=1-2x Now that we know what y=, we ca fit it into the second equation. 4x+2(1-2x)=-1 Now we disribute 4x+2-4x=-1 We have like terms, so 4x-4x=0x Now we are left wsith 2=-1 This is not a solution, so no solution.

OpenStudy (cedric):

hmmm i like the wway you worded that, can yall help with one more?

OpenStudy (anonymous):

Sure, if it's an equation.

OpenStudy (cedric):

Stable High School has a total of 112 boys and girls who play sports. If the number of girls, g is 16 more than twice the number of boys, b, how many girls play sports at this high school? Which system of equations would help you solve this problem? b + g = 112 b = 2g + 16 b + g = 112 g = 2b + 16 b − g = 112 b = 2g − 16 None of the systems above would solve this problem

OpenStudy (anonymous):

Please post a new question in a new post...

OpenStudy (cedric):

oh ok